If an 1800-kcalorie diet contains 100 grams of fat, the percentage of kcal from the fat is?
Answer #1

ANSWER.

* Calculation of percentage kcal from fat.

         Total kcal in diet = 1800 kcal.

         Total fat (gm) in the diet = 100 gm

               1 gm fat = 9 kcal.

     So

        Total fat (kcal) in the diet = 100 * 9

                                                = 900 kcal.

   Percentage of kcal from the fat = Total kcal of fat in the diet / Total kcal * 100

                                                      = 900 / 1800 * 100

                                                      = 0.5 * 100

                                                      = 50 %